OpenStudy (anonymous):

Can you solve the equation 2/3(4x-5)=8 by using the Division Property of Equality? Explain.

OpenStudy (alienium):

2/3 * (4x-5) = 8 idk what that is but i can solve it: (8x-10)/3=8 8x-10=24 8x=24+10 x=34/10 x=3.4

OpenStudy (anonymous):

x is not 3.4

OpenStudy (anonymous):

x is 17/4

OpenStudy (anonymous):

I just need an explanation to why You can use the Division Property of Equality with this equation

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }(4x-5)=8\] ????

OpenStudy (alienium):

just because the question says so... you can solve it pretty much with 2 moves

OpenStudy (anonymous):

ok